Fill in the gaps in the table below 


QuantityofVariableinputTotalOutputMarginalProductofVariableinputAverageProduut

 0 0

1 225

2 300

3 300

4 1140

5 225

6 225


1
Expert's answer
2021-12-22T09:24:49-0500

Total output of input 2 = Average product × 2 = 300 × 2 = 600

Total output of input 6 = Average product × 6 = 225 × 6 = 1,350

Marginal product input 1 = Total output of input 1 – Total output 0 = 225 - 0 = 225

Marginal product input 2 = Total output of input 2 – Total output 1 = 600 - 225 = 375

Marginal product input 4 = Total output of input 4 – Total output 3 = 1140 - 900 = 240

Marginal product input 6 = Total output of input 6 – Total output 5 = 1350 - 1365 = -15


The complete table is as below: